OpenStudy (lexi724):

MEDAL FOR CHECKING/Helping:

OpenStudy (lexi724):

find the side lengths

OpenStudy (anonymous):

Well i Officially Feel not qualified for ANYTHING! sorry i dont know how to help you with that.

OpenStudy (lexi724):

:-( tis okay doll

OpenStudy (anonymous):

but if it helps heres a medal

OpenStudy (lexi724):

haha thanks XD

OpenStudy (lexi724):

maybe @ankitshaw can help

OpenStudy (anonymous):

exactly waht i was gonna say

OpenStudy (ankitshaw):

what is your answer?

OpenStudy (lexi724):

i dont know how to do this the checking is by making sure im solving correctly as im getting help

OpenStudy (lexi724):

@Ziferot knows how to help for sure im just probably annoying him tho

OpenStudy (ankitshaw):

ok so see in the triangle hypotenuse is given as 30 .. and we have angle EFD 30.. use sin = side/hypotenuse to find the side lenght ED

OpenStudy (ankitshaw):

sorry hypotenuse is 40

OpenStudy (lexi724):

sin=60*40?

OpenStudy (lexi724):

or -0.3048106211 *40

OpenStudy (ankitshaw):

sin30= lenght ED/EF

OpenStudy (lexi724):

-0.98803162409*60*30

OpenStudy (ankitshaw):

\[\sin 30= \frac{ ED }{ Ef } \] \[\[\sin 30= \frac{ ED }{ 40 } \]\] \[\[\ \frac{ 1 }{ 2 }= \frac{ ED }{ 40 } \]\]

OpenStudy (lexi724):

1/2*60 divided by 40?

OpenStudy (lexi724):

1.5*1/2=0.75

OpenStudy (lexi724):

@ankitshaw ...........

OpenStudy (ankitshaw):

\[\frac{ 1 }{ 2 } = \frac{ ED }{ 40 }\] so when you take 40 on the other side you get 40/2 ..that is 20 so ED is 20

OpenStudy (lexi724):

ok so 20*40=

OpenStudy (lexi724):

cuz side*hypotnuse

OpenStudy (ankitshaw):

what is side*hypotnuse ?

OpenStudy (lexi724):

thats what you said earlier

OpenStudy (ankitshaw):

no i said side/hypotenuse not side*hypotenuse

OpenStudy (lexi724):

20/40 then

OpenStudy (lexi724):

0.5

OpenStudy (ankitshaw):

no no.. sin30 = side / hypotenuse was the formula used to find the side lenght ED

OpenStudy (lexi724):

i give up

OpenStudy (ankitshaw):



OpenStudy (lexi724):

ok

OpenStudy (lexi724):

my class is going to be way over before i figure the rest of these out and i have 7 more

OpenStudy (anonymous):

i have some work, u need help?

OpenStudy (lexi724):

yes @Ziferot

OpenStudy (anonymous):

whats the question??

OpenStudy (lexi724):

find the side lengths

OpenStudy (anonymous):

40 * sin 30 40 * sin 60

OpenStudy (anonymous):

or 40 * sin30 40 * cos 30

OpenStudy (lexi724):

-39.5 6.17

OpenStudy (lexi724):

@Ziferot

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

sin 30 = 1/2

OpenStudy (anonymous):

sin 60 = sqrt3/2

OpenStudy (lexi724):

1/2*30=15

OpenStudy (anonymous):

u have 40

OpenStudy (anonymous):

not 30

OpenStudy (lexi724):

i give up im just so done thank you guys but im retarded so i cant do this

OpenStudy (anonymous):

it will be A